Find
dz/dt
for z = x^2 y + 4y^2 where x = cost and y = sin t using the chain rule
1
Expert's answer
2020-04-06T16:37:57-0400

Given that "z=x^2y+4y^2" and "x=cost \ , \ y=sint" .



"\therefore \frac{dz}{dt}=\frac{d(x^2y+4y^2)}{dt}"

"\implies \frac{dz}{dt}=\frac{d(x^2y)}{dt} +\frac{d(4y^2)}{dt}""\implies \ \frac{dz}{dt}=x^2\frac{dy}{dt}+y\frac{dx^2}{dt}+4\frac{dy^2}{dt}"


Now ,using chain rule we get


"\frac{dz}{dt}=x^2\frac{dy}{dt}+y\frac{dx^2}{dx} ×\frac{dx}{dt}+4\frac{dy^2}{dy}×\frac{dy}{dt}"


Since,

"x=cost \ \implies \frac{dx}{dt}=-sint""and \ y=sint \ \implies \ \frac{dy}{dt}=cost""\therefore \ \frac{dz}{dt}=x^2cost-2xysint+8ycost."